Kingbird Inc., a private company following ASPE, is having difficulty meeting its working capital requirements. As a result, on January 1, 2020, the company sold bonds with a face value of $1.05 million, receiving $850,000 in cash. The bonds have an interest rate of 6% and mature on January 1, 2022. Interest is payable semi-annually on January 1 and July 1.

Set up a schedule of interest expense and discount amortization under the straight-line method.
Answer #1

A B C D E F 1 Date Credit interest payable Debit interest expense Credit discount bond payable (Amortization) Jan. 1, 2020 Ju

Working note:

B D Credit Debit Date interest interest payable expense Jan. 1, 2020 July. 1, 2020 =1050000*6% =C4+E4 Jan. 1, 2021 =1050000*6